# Beyond 1D: Vision Transformers and Multichannel Signal Images for PPG-to-ECG Reconstruction

### v1

@@ 7 Conclusion
This study presents a novel four-channel image-based approach for ECG reconstruction from PPG, leveraging a ViT architecture. Unlike conventional 1D convolution-based methods, our approach utilizes a four-channel image representation that retains richer PPG information and effectively captures both intra-beat features and inter-beat relationships.
Experimental results show that our method consistently outperforms state-of-the-art 1D convolution-based models across almost all evaluation metrics, achieving up to 29% reduction in PRD and 15% reduction in RMSE. To ensure proper beat alignment, we padded shorter beats with subsequent beats instead of zero values, preserving accurate patch representations at beat transitions and maintaining beat-to-beat continuity. Additionally, comparisons between the ViT and the Transformer highlight ViT’s superior ability to model beat-to-beat variations and capture complex temporal dependencies.
Despite these advantages, certain design choices impact performance, particularly image size and patch size selection. Given that we use two datasets, we set each signal image to include 16 padded beats with a patch size of $8\times 8$. Theoretically, increasing the number of beats per image and using a larger patch size could enable the model to better capture beat-to-beat variations. However, this approach reduces the number of training samples, which can lead to data scarcity issues and decreased model generalization.
Beyond advancing non-invasive cardiovascular monitoring, our approach introduces a new paradigm for PPG feature extraction and opens new avenues for cyclic signal prediction. However, accurately identifying the onset and offset of small waveform features, such as P-wave onset and T-wave offset, remains challenging due to noise interference. To mitigate this, we approximated intervals like the PR interval using peak distances instead of onset markers. While this provides a reliable foundation for error analysis, future research should focus on improving small waveform boundary detection to enhance clinical applicability.
Appendix A Additional ECG Signal Reconstructions
Figures 12 and 13 present additional visualizations of reconstructed ECG signals from two datasets. These examples demonstrate that our proposed method achieves promising results in PPG-to-ECG reconstruction, highlighting its potential for using PPG as an alternative measurement of heart activity.
@@ Appendix A Additional ECG Signal Reconstructions
Appendix B Visualization of Four-Channel Output
(a) Generated four-channel ECG image.
Figure 14: Visualization of the four-channel output and corresponding signals for record 0325 in the CapnoBase dataset. The output is generated using the weighted sum-based combination approach with two reconstruction losses ($L_{R_{1C}}$ + $L_{R_{4C}}$). Figures 14(a) and 14(b) depict the intermediate four-channel ECG image and its corresponding signal, while Figure 14(c) illustrates the final reconstructed ECG signal after padding removal.

### v2

@@ 7 Conclusion
This study presents a novel approach for ECG reconstruction from PPG by leveraging a ViT and a four-channel image-based signal representation. Unlike traditional 1D CNN-based models, our method reformulates PPG sequences into structured 2D signal images comprising the raw waveform, its first- and second-order differences, and the AUC. This enriched multi-channel representation enables the ViT to model both intra-beat morphology and cross-cycle temporal dynamics through self-attention mechanisms.
The proposed framework demonstrates strong empirical performance, achieving up to a 29% reduction in PRD and a 15% reduction in RMSE compared to state-of-the-art baselines. To preserve temporal continuity across cardiac cycles, we introduced a beat-padding strategy that maintains beat-to-beat continuity better than conventional zero-padding. Additionally, we employed clinically relevant metrics, such as QRS area and RT interval errors, to enhance interpretability beyond conventional point-wise evaluations.
Despite its strengths, the current implementation has limitations. Chief among them is the use of a fixed patch size in the ViT architecture, which may not generalize well to signals with varying sampling rates or beat durations. In high-frequency signals, large patches may span multiple waveform components, reducing the model’s ability to resolve distinct features. Moreover, static patching may miss subtle yet clinically important morphological variations, particularly in datasets with high inter-subject variability.
Nonetheless, our proposed framework provides a solid foundation for developing practical algorithms for ECG reconstruction. Future work should explore adaptive or learnable patching strategies that dynamically align patch size with signal properties. Integrating auxiliary networks trained to detect specific waveform boundaries, such as the onset of the P wave or the offset of the T wave, may enhance boundary localization and improve clinical interpretability. Furthermore, extending this approach to multi-lead ECG reconstruction and evaluating its robustness under motion artifacts or across wearable PPG devices would enhance its real-world applicability. Beyond ECG reconstruction, the proposed ViT-based framework provides a generalizable solution for modeling cyclic physiological signals, such as respiratory or arterial waveforms, contributing broadly to biomedical signal processing and pattern recognition.
Appendix A Visualization of Four-Channel Output
(a) Generated four-channel ECG image.
Figure 12: Visualization of the four-channel output and corresponding signals for record 0325 in the CapnoBase dataset. The output is generated using the weighted sum-based combination approach with two reconstruction losses ($L_{R_{1C}}$ + $L_{R_{4C}}$). Figures 12(a) and 12(b) depict the intermediate four-channel ECG image and its corresponding signal, while Figure 12(c) illustrates the final reconstructed ECG signal after padding removal.
Appendix B Additional ECG Signal Reconstructions
Figures 13 and 14 present additional visualizations of reconstructed ECG signals from two datasets. These examples demonstrate that our proposed method achieves promising results in PPG-to-ECG reconstruction, highlighting its potential for using PPG as an alternative measurement of heart activity.